Consider the matrix A="\begin{bmatrix}
 1 & 0 & 2 \\
 0 & 1 & 0 \\
 2 & 0 & 1
\end{bmatrix}"

1.1. Show that the characteristic equation for the eigenvalues "\lambda" of A is given by (λ2-1) (λ-3)=0.

1.2. Find an orthogonal matrix P which diagonalizes A.

1.3. Find An (for n ∈ N) as a matrix


1
Expert's answer
2021-11-02T15:23:53-0400

"|A-\lambda I|=\begin{vmatrix}
 1-\lambda & 0&2 \\
 0& 1-\lambda&0\\
2&0&1-\lambda
\end{vmatrix}=0\\
(1-\lambda)^3-0-0-4(1-\lambda)-0-0=0\\
(1-\lambda)((1-\lambda)^2-4)=0\\
(1-\lambda )(1-\lambda-2)(1-\lambda+2)=0\\
(1-\lambda)(-1-\lambda)(3-\lambda)=0\\
(\lambda^2-1)(\lambda-3)=0\\
\lambda_1=1, \lambda_2=-1, \lambda_3=3"

"\lambda_1=1, \vec{a_1}=(x,y,z)\neq\vec{0}\\
(A-\lambda_1I)\vec{a_1}=\vec{0}\\
\begin{pmatrix}
 0& 0&2 \\
 0 & 0&0\\
2&0&0
\end{pmatrix}\cdot \begin{pmatrix}
 x \\
 y\\z
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0\\0
\end{pmatrix}\\
2z=0,2x=0\\
x=z=0, y\in R\\
\vec{a_1}=(0,1,0)"

"\lambda_2=-1, \vec{a_2}=(x,y,z)\neq\vec{0}\\
(A-\lambda_2I)\vec{a_2}=\vec{0}\\
\begin{pmatrix}
 2& 0&2 \\
 0 & 2&0\\
2&0&2
\end{pmatrix}\cdot \begin{pmatrix}
 x \\
 y\\z
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0\\0
\end{pmatrix}\\
2x+2z=0,2y=0\\
x=-z, y=0\\
\vec{a_2}=(1,0,-1)"


"\lambda_3=3, \vec{a_3}=(x,y,z)\neq\vec{0}\\
(A-\lambda_3I)\vec{a_3}=\vec{0}\\
\begin{pmatrix}
 -2& 0&2 \\
 0 & -2&0\\
2&0&-2
\end{pmatrix}\cdot \begin{pmatrix}
 x \\
 y\\z
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0\\0
\end{pmatrix}\\
-2x+2z=0,-2y=0, 2x-2z=0\\
x=z=1, y=0\\
\vec{a_3}=(1,0,1)"

"\vec{a_1}=(0,1,0), |\vec{a_1}|=1,\\
\vec{a_2}=(1,0,-1), |\vec{a_2}|=\sqrt2,\\
\vec{a_3}=(1,0,1),|\vec{a_3}|=\sqrt2\\
P=\begin{pmatrix}
 0 & 1&0 \\
 \frac{1}{\sqrt2} & 0& -\frac{1}{\sqrt2}\\
 \frac{1}{\sqrt2}&0 &\frac{1}{\sqrt2}
\end{pmatrix}\\
B=P^{-1}AP\\
B=\begin{pmatrix}
 1 & 0&0 \\
 0 & -1&0\\
0&0&3
\end{pmatrix}"


"A^2=\begin{pmatrix}
 1 & 0&2 \\
 0&1&0\\
2&0&1
\end{pmatrix}\cdot\begin{pmatrix}
 1 & 0&2 \\
 0&1&0\\
2&0&1
\end{pmatrix}=\begin{pmatrix}
 5 & 0&4 \\
 0&1&0\\
4&0&5
\end{pmatrix}\\
A^3=A^2\cdot A=\begin{pmatrix}
 5 & 0&4 \\
 0&1&0\\
4&0&5
\end{pmatrix}\cdot \begin{pmatrix}
 1 & 0&2 \\
 0&1&0\\
2&0&1
\end{pmatrix}=\\
=\begin{pmatrix}
 13 & 0&14 \\
 0&1&0\\
14&0&13
\end{pmatrix}"


"A^4=A^3\cdot A=\begin{pmatrix}
 13 & 0&14 \\
 0&1&0\\
14&0&13
\end{pmatrix}\cdot \begin{pmatrix}
 1 & 0&2 \\
 0&1&0\\
2&0&1
\end{pmatrix}=\\
=\begin{pmatrix}
 41 & 0&40 \\
 0&1&0\\
40&0&41
\end{pmatrix}"


"A^n=\begin{pmatrix}
 a & 0&b \\
 0&1&0\\
b&0&a
\end{pmatrix}"